OpenStudy (anonymous):

Determine wether the expression is a polynomial. If so how many terms and variables does the polynomial have. Then state its degree. 5x-1/7x (the 1/7 is a fraction and the 7 has the x) Help Please

OpenStudy (anonymous):

@jim_thompson5910 can you please help me? sorry to be such a bother

jimthompson5910 (jim_thompson5910):

is the expression \[\Large \frac{5x - 1}{7x}\] or is it \[\Large \frac{5x - 1}{7}x\] or is it \[\Large 5x - \frac{1}{7}x\]

OpenStudy (anonymous):

The last one @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok so because you're not dividing by any variables, it's a polynomial the degree is equal to the largest exponent

OpenStudy (anonymous):

how many terms and variables would it have? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

a term is basically a monomial a monomial is something of the form (coefficient)*(variable)

jimthompson5910 (jim_thompson5910):

they are separated by a + or a -

jimthompson5910 (jim_thompson5910):

ex: 2x + 5 has two terms: 2x and 5

OpenStudy (anonymous):

so the terms are 5x and the fraction \[\frac{ 1 }{ 7x }\] what about the variables? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the variables are just the unknowns, so in this case, just x

OpenStudy (anonymous):

ok thank you and what did you mean by the highest exponent? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well if you had something like 3x^7 + 9x^5 - 10x^3 - 4 then the largest exponent is 7 so the degree of that is 7

OpenStudy (anonymous):

in this case would it be 0 or 1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

5x is the same as 5x^1 (1/7)x is the same as (1/7)x^1

OpenStudy (anonymous):

ok then 1 is the degree @jim_thompson5910 Thank You Very Much

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

one more please? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

4x^-3y^2 same question different equation @jim_thompson5910

OpenStudy (anonymous):

so this is a polynomial correct? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

-3 is the exponent?

OpenStudy (anonymous):

yes @jim_thompson5910 and 2 is the exponent for y

jimthompson5910 (jim_thompson5910):

ok so because you have a negative exponent, it cannot be a polynomial

jimthompson5910 (jim_thompson5910):

negative exponents ---> not a polynomial

OpenStudy (anonymous):

ok thank you again @jim_thompson5910 && can you check my work on some of my questions please?

OpenStudy (anonymous):

sorry to be such a bother

jimthompson5910 (jim_thompson5910):

sure i can do a few

OpenStudy (anonymous):

\[(2a ^{3}b ^{2}+6a ^{4}b) + (a ^{4}b -2a ^{3}b ^{2})=2a ^{7}b ^{3}+ 4a ^{1}b\] is that answer correct? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no it is not

OpenStudy (anonymous):

what did i do wrong? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

2a^3b^2 + (-2a^3b^2) = 2a^3b^2 - 2a^3b^2 = 0a^3b^2 = 0 so those two cancel out

jimthompson5910 (jim_thompson5910):

6a^4b + a^4b = 7a^4b

jimthompson5910 (jim_thompson5910):

So the answer is really \[\Large 7a^{4}b\]

OpenStudy (anonymous):

oh ok i understand now for this one......... (11-z)z = 11z-z^2 @jim_thompson5910 is that right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

it's correct

OpenStudy (anonymous):

thank you...... 6xy(x+y) = 6x^2+6xy^2 is that the answer? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it should be 6x^2y+6xy^2 notice how there is a y term after the 6x^2

OpenStudy (anonymous):

oh yes sorry forgot to put that thanks and (4v-1)(5v-3)= 20v^2-17v+3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

Thanks=) ....... (4v+4)(4v-4)=16v^2-16 @jim_thompson5910

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

Thank you again and just two more please? sorry to be a bother i just need to do good @jim_thompson5910

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

yay thank you so much (x^2+y^2)(x^2-y^2)= x^4-y^4 is that right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes it is, you're on a roll

OpenStudy (anonymous):

OK this one im not so sure of.... \[\frac{ x ^{8} +x ^{2} }{ x } = x ^{7}x\] would that be right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

x^7 + x, but you were close

jimthompson5910 (jim_thompson5910):

\[\Large x^{7}+x\]

OpenStudy (anonymous):

I cannot tell you how grateful i am thank you very very much you really are an amazing tutor @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

do i have to put any parenthisis for any of the answers or no? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

yes for like the ones with multyplying @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no you don't need to put parenthesis in the answers

jimthompson5910 (jim_thompson5910):

something like (x+5) is the same as x+5

jimthompson5910 (jim_thompson5910):

when you factor, you need parenthesis, but I don't think you factored any of them

OpenStudy (anonymous):

ok thank you again @jim_thompson5910 hopefully i get 100% @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm sure you will

OpenStudy (anonymous):

i just got one wrong @jim_thompson5910

jimthompson5910 (jim_thompson5910):

which one was that

OpenStudy (anonymous):

number 1 the first question it said it wasnt a polynomial @jim_thompson5910

jimthompson5910 (jim_thompson5910):

hmm i guess x was in the denominator

jimthompson5910 (jim_thompson5910):

or something was missing there

OpenStudy (anonymous):

i can redo it and get 100 on it can you help again? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have to do the whole thing over?

OpenStudy (anonymous):

yes different question thereis only 9 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok not so bad i guess

OpenStudy (anonymous):

thank you=) first question>>>>> \[5x+\frac{ 1 }{ 3x }\] this is a polynomial the polynomial has 2 terms and 2 variables the highest degree is 1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no it is not

jimthompson5910 (jim_thompson5910):

you're dividing by a variable, so that's why its not

jimthompson5910 (jim_thompson5910):

that's probably why we missed the first one btw i think you confused one of the expressions above and I used the wrong formula, but it's ok

OpenStudy (anonymous):

oh ok sorry.... \[6x ^{-1}y ^{-2}\] not a polynomial @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

correct, \[6x ^{-1}y ^{-2}\] is not a polynomial

OpenStudy (anonymous):

(a^3b^2+5ab^4)+(ab^4-a^3b^2)= 5a^7b^8 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no that's not correct

jimthompson5910 (jim_thompson5910):

5ab^4 + ab^4 = ??

jimthompson5910 (jim_thompson5910):

you do NOT add the exponents btw

jimthompson5910 (jim_thompson5910):

just the coefficients for the like terms

OpenStudy (anonymous):

6ab^4? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's your answer

jimthompson5910 (jim_thompson5910):

the other pair of terms cancel out and go away

OpenStudy (anonymous):

so \[6ab ^{4}\] is the answer? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

(6-z)z = 6z-z^2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

12xy(x+y)= 12x^2y + 12xy^2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

(4r-1)(5r-3)= 20r^2-17r+3 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you got it right

OpenStudy (anonymous):

(2b+6)(2b-6)= 4b^2-36 @jim_thompson5910